# Consistent Validation for Predictive Methods in Spatial Settings

### v1

@@ Appendix B Proofs of Claims
We now present results and proofs not included in the main text. We essentially follow the order of results in the main text. Section Section B.1 gives sufficient conditions for Assumption 2.5 to hold for a homoskedastic, additive noise model and squared loss for responses taking values in $[0,1]$, mentioned in Section 2.2. Section B.2 focuses on the proof of Prop. 3.3, first recalling several properties of covering numbers that will be used in the result. In Section B.3 we restate and prove our Prop. 4.1 and Prop. 4.3, which show limitations of existing validation methods in the spatial setting we study. In Section B.4 we prove a general result upper bounding the error in estimating the risk using $k$-nearest neighbors, as well as an upper bound on the error for $k^{\star}_{T_{2}}$. Section B.5 establishes the consistency of spatial nearest neighbors. Section C.2 discusses issues related to model selection and proves rates of convergence for spatial nearest neighbors.
B.1 Lipschitz Constant for Lipschitz Response and Predictive Method
@@
Proposition B.1.
Consider $\mathcal{Y}\subset[0,1]$ and squared loss. Then if $f^{\chi}(S):=f(S,\chi(S))$ is $L_{Y}$-Lipschitz and $h^{\chi}$ is $L_{h}$-Lipschitz. Let $(S,X,Y)$ be generated as in Assumption 2.2, with $Y=f(S,X)+\epsilon$. Then $e^{\chi}_{h}(S)\coloneqq\mathbb{E}[(Y-h^{\chi}(S))^{2}|S]$ is $2(L_{Y}+L_{h})$-Lipschitz.
@@
$|e^{\chi}_{h}(S)-e^{\chi}_{h}(S^{\prime})|$ $=\mathbb{E}[(f^{\chi}(S)+\epsilon)-h^{\chi}(S))^{2}-(f^{\chi}(S^{\prime})+\epsilon^{\prime})-h^{\chi}(S^{\prime}))^{2}]$ (17)
@@
$f^{\chi}(S)^{2}-f^{\chi}(S^{\prime})^{2}+h^{\chi}(S)^{2}-h^{\chi}(S^{\prime})^{2}$ $=(f^{\chi}(S)+f^{\chi}(S^{\prime}))(f^{\chi}(S)-f^{\chi}(S^{\prime}))+(h^{\chi}(S)+h^{\chi}(S^{\prime}))(h^{\chi}(S)-h^{\chi}(S^{\prime}))$
@@
Therefore, at least in the case of squared loss with bounded response and predictive method values and a homoskedastic additive noise model, smoothness of the average response surface (as a function of space) together with smoothness of the predictive method imply Assumption 2.5.
B.2 Proof that Independent and Identically Distributed Data Implies Infill Asymptotic with High Probability
The purpose of this section is to prove Prop. 3.3. We begin by recalling this proposition:
@@
Our proof is similar to earlier proofs in Reznikov & Saff (2015, Section 5.2) and essentially follows the stackexchange response (2021) (https://mathoverflow.net/users/36721/iosif pinelis) keeping track of numerical constants. Essentially, the idea is that were the fill distance to be large, there must be a ball of large radius that doesn’t contain any points in the sample of location. But because the probability of a point in the sample falling in any ball is bounded below in terms of the volume of the ball, this must be improbable as the sample size grows.
B.2.1 Preliminary Definitions and Lemmas Related to Nets and Covering Number
In order to formalize the proof of Prop. 3.3 sketched above idea, we recall the definition of a net and covering number, as well as some standard properties of covering numbers.
Definition B.2 (Net, Covering Number).
Let $A\subset\mathbb{R}^{d}$ compact. Any finite set $C\subset\mathbb{R}^{d}$ such that
@@ B.2.1 Preliminary Definitions and Lemmas Related to Nets and Covering Number
We now recall how covering number behaves under affine transformations of the underlying space. For set $A,B$, we define $A+B=\{a+b:a\in A,b\in B\}$.
Proposition B.3.
@@
In the proof of Prop. 3.3 we apply a union bound over all elements in a net for the unit cube. We therefore need a result telling us that this net does not contain too many elements.
Lemma B.4 (Covering number of Unit Cube).
@@
By the upper bound in Wainwright (2019, Lemma 5.7),
@@
B.2.2 Main Proof
We recall and prove Prop. 3.3.
The proof essentially follows (2021) (https://mathoverflow.net/users/36721/iosif pinelis), making constants explicit. The idea of the proof is that the distance from any point in $[0,1]^{d}$ can’t be larger than the distance from this point to the nearest point in some net, plus the distance from that point to the nearest point in the validation sample. Therefore, it suffices to show that at least one point in the sample falls in a ball of some radius around each point in a net. This is done with a union bound over all elements in a minimal net, using that the sampling density of validation points is bounded below and so the probability of any particular validation point falling in a ball centered at any point in the net can’t be too small.
We again recall Prop. 3.3 for convenience when reading the proof. See 3.3
Proof of Prop. 3.3.
Fix $\tau\in(0,1)$ and let $C$ be a minimal cardinality $\tau/2$-net for $[0,1]^{d}$. If $B(S,\tau/2)$ contains a validation point for all $S\in C$, then for any $S^{\prime}\in[0,1]^{d}$, by the triangle inequality,
@@ B.2.2 Main Proof
Returning to Eqn. 27 and taking a union over all elements in the net:
@@
where $B_{d}=\frac{\pi^{d/2}}{\Gamma(\frac{d}{2}+1)}$ is the volume of the $d$-dimensional unit sphere. Eqn. 30 uses Eqn. 28. We now upper bound the terms in Eqn. 31.
Applying, Props. B.3 and B.4
@@
Now choose $\tau^{d}=\frac{4^{d}}{cN^{\textup{{val}}}B_{d}}\log\frac{6^{d}N^{\textup{{val}}}}{B_{d}\delta}$. For all $N^{\textup{{val}}}$ larger than some $n_{0}$, $\tau<1$ because $\lim_{N^{\textup{{val}}}\to\infty}\frac{4^{d}}{cN^{\textup{{val}}}B_{d}}\log\frac{6^{d}N^{\textup{{val}}}}{B_{d}\delta}=0$. For this choice of $\tau$
@@
where in the final inequality we use that $B_{d}\leq 2^{d}$ since the unit ball is contained in the unit cube for $\tau<1$. For all $N^{\textup{{val}}}\geq\frac{\delta e^{c/2^{d}}}{3^{d}}$ the right-hand side is less than $\delta$; but $\delta<1,c\leq 1$ and so this holds for all $N^{\textup{{val}}}\geq 1$. ∎
B.3 Proof of Inconsistency of Existing Methods
@@ B.3.1 Inconsistency of the Holdout
See 4.1
@@
Because $e^{\chi}_{h}(S)$ is not a constant $\mathcal{S}$ contains at least two elements $S,S^{\prime}$ such that $e^{\chi}_{h}(S)\neq e^{\chi}_{h}(S^{\prime})$. Let $Q=\delta_{S}$ and $Q^{\prime}=\delta_{S^{\prime}}$. There there exists some $\gamma>0$ such that
$|R^{\chi}_{Q}(h)-R^{\chi}_{Q^{\prime}}(h)|=|e^{\chi}_{h}(S)-e^{\chi}_{h}(S^{\prime})|>2\gamma.$ (36)
If $\hat{R}^{\chi}_{\textup{{Hold}}}$ is not consistent for $R^{\chi}_{Q}(h)$, we are done. Otherwise, by the definition of consistency, for all $N\geq N_{0}$ with probability at least $1/2$,
$|(\hat{R}^{\chi}_{\textup{{Hold}}})^{(N)}(h)-R^{\chi}_{Q}(h)|<\gamma.$ (37)
where we use $(\hat{R}^{\chi}_{\textup{{Hold}}})^{(N)}$ to denote the estimator constructed using the first $N$ validation points, $D^{\textup{{val}}}_{N}$. By the reverse triangle inequality,
$|(\hat{R}^{\chi}_{\textup{{Hold}}})^{(N)}(h)-R^{\chi}_{Q^{\prime}}(h)|\geq|R^{\chi}_{Q}(h)-R^{\chi}_{Q^{\prime}}(h)|-|(\hat{R}^{\chi}_{\textup{{Hold}}})^{(N)}(h)-R^{\chi}_{Q^{\prime}}(h)|.$ (38)
Combining Eqn. 37 and Eqn. 38, for all $N\geq N_{0}$ with probability at least $1/2$
$|(\hat{R}^{\chi}_{\textup{{Hold}}})^{(N)}(h)-R^{\chi}_{Q^{\prime}}(h)|>2\gamma-\gamma=\gamma,$ (39)
which implies the holdout is not consistent for $R^{\chi}_{Q^{\prime}}(h)$. ∎
B.3.2 Inconsistency of $1$-nearest neighbor Estimator
We now turn to $1$-nearest neighbor risk estimation and restate and prove Prop. 4.3: See 4.3
@@ B.3.2 Inconsistency of $1$-nearest neighbor Estimator
B.3.3 Preliminary Result
We begin by proving a result that says that if a bounded random variable has a second moment bounded below by $C$, then it cannot be close to zero most of the time. We will apply this inequality to the second moment of the difference between the $1$-nearest neighbor estimator and the test risk to in our proof of Prop. 4.3.
Proposition B.5.
@@ B.3.3 Preliminary Result
Because $U\in[-A,A]$, $A^{2}-U^{2}$ is a non-negative random variable. Applying Markov’s inequality, for any $t\geq 0$,
@@
Take $t=\frac{A^{2}-C}{\delta}$ which is greater than 00 because $\delta\in(1-\frac{C}{A^{2}},1)$. Then Eqn. 41 becomes,
@@
B.3.4 Restatement and Proof of Inconsistency of $1$-nearest neighbor
We return to our proof of Prop. 4.3. The idea will be to consider a point prediction task and then apply Prop. B.5 to show that with some fixed probability, the $1$-nearest neighbor estimator is a fixed distance away from the test risk, even as $N^{\textup{{val}}}$ increases.
See 4.3

### v2

@@ Appendix D Proofs of Claims
We now present results and proofs not included in the main text. We essentially follow the order of results in the main text. Section Section D.1 gives sufficient conditions for Assumption 2.5 to hold for a homoskedastic, additive noise model and squared loss for responses taking values in $[0,1]$, mentioned in Section 2.2. Section D.2 focuses on the proof of Proposition 3.3, first recalling several properties of covering numbers that will be used in the result. In Section D.3 we restate and prove our Propositions 4.1, 4.3 and 4.4, which show limitations of existing validation methods in the spatial setting we study. In Section D.4 we prove a general result upper bounding the error in estimating the risk using $k$-nearest neighbors, as well as an upper bound on the error for $k^{\star}_{T_{2}}$. Section D.5 establishes the consistency of spatial nearest neighbors. Section D.6 discusses issues related to model selection and proves rates of convergence for spatial nearest neighbors.
D.1 Lipschitz Constant for Lipschitz Response and Predictive Method
@@
Proposition D.1.
Consider $\mathcal{Y}\subset[0,1]$ and squared loss. Suppose $f^{\chi}(S):=f(S,\chi(S))$ is $L_{Y}$-Lipschitz and $h^{\chi}$ is $L_{h}$-Lipschitz. Let $(S,X,Y)$ be generated as in Assumption 2.2, with $Y=f(S,X)+\epsilon$. Then $e_{h}(S)\coloneqq\mathbb{E}[(Y-h^{\chi}(S))^{2}|S]$ is $2(L_{Y}+L_{h})$-Lipschitz.
@@
$|e_{h}(S)-e_{h}(S^{\prime})|$ $=\mathbb{E}[(f^{\chi}(S)+\epsilon)-h^{\chi}(S))^{2}-(f^{\chi}(S^{\prime})+\epsilon^{\prime})-h^{\chi}(S^{\prime}))^{2}]$ (5)
@@
$f^{\chi}(S)^{2}-f^{\chi}(S^{\prime})^{2}+h^{\chi}(S)^{2}-h^{\chi}(S^{\prime})^{2}$ $=(f^{\chi}(S)+f^{\chi}(S^{\prime}))(f^{\chi}(S)-f^{\chi}(S^{\prime}))$ (7)
$\phantom{\qquad}+(h^{\chi}(S)+h^{\chi}(S^{\prime}))(h^{\chi}(S)-h^{\chi}(S^{\prime}))$
@@
Therefore, at least in the case of squared loss with bounded response and predictive method values and a homoskedastic additive noise model, smoothness of the average response surface (as a function of space) together with smoothness of the predictive method imply Assumption 2.5. We expect to hold for other losses that are Lipschitz functions of the response and prediction.
D.2 Proof that Independent and Identically Distributed Data Implies Infill Asymptotic with High Probability
The purpose of this section is to prove Proposition 3.3. We begin by recalling this proposition:
@@
Our proof is similar to earlier proofs in Reznikov and Saff [2015, Section 5.2] and essentially follows the stackoverflow response (2021) [https://mathoverflow.net/users/36721/iosif pinelis] keeping track of numerical constants. Essentially, the idea is that were the fill distance to be large, there must be a ball of large radius that doesn’t contain any points in the sample of location. But because the probability of a point in the sample falling in any ball is bounded below in terms of the volume of the ball, this must be improbable as the sample size grows.
D.2.1 Preliminary Definitions and Lemmas Related to Nets and Covering Number
In order to formalize the proof of Proposition 3.3 sketched above idea, we recall the definition of a net and covering number, as well as some standard properties of covering numbers.
Definition D.2 (Net, Covering Number).
Let $A\subset\mathbb{R}^{d}$ be a compact set. Any finite set $C\subset\mathbb{R}^{d}$ such that
@@ D.2.1 Preliminary Definitions and Lemmas Related to Nets and Covering Number
We now recall how covering number behaves under affine transformations of the underlying space. For sets $A,B$ and a scalar $\alpha$, we define $A+B=\{a+b:a\in A,b\in B\}$ and $\alpha A=\{\alpha a:a\in A\}$.
Proposition D.3.
@@
In the proof of Proposition 3.3 we apply a union bound over all elements in a net for the unit cube. We therefore need a result telling us that this net does not contain too many elements.
Lemma D.4 (Covering number of Unit Cube).
@@
By the upper bound in Wainwright [2019, Lemma 5.7],
@@
D.2.2 Main Proof
We again recall Proposition 3.3 for convenience when reading the proof. See 3.3
Proof of Proposition 3.3.
For some $\tau\in(0,1)$ (to be selected later) let $C$ be a minimal cardinality $\tau/2$-net for $[0,1]^{d}$. If $B(S,\tau/2)$ contains a validation point for all $S\in C$, then for any $S^{\prime}\in[0,1]^{d}$, by the triangle inequality,
@@ D.2.2 Main Proof
Returning to Equation 16 and taking a union over all elements in the net:
@@
where $B_{d}=\frac{\pi^{d/2}}{\Gamma(\frac{d}{2}+1)}$ is the volume of the $d$-dimensional unit sphere. Equation 19 uses Equation 17. We now upper bound the terms in Equation 20.
Applying, Propositions D.3 and D.4
@@
Now choose $\tau^{d}=\frac{4^{d}}{cN^{\textup{{val}}}B_{d}}\log\frac{6^{d}N^{\textup{{val}}}}{B_{d}\delta}$. For all $N^{\textup{{val}}}$ larger than some $n_{0}$, $\tau<1$ because $\lim_{N^{\textup{{val}}}\to\infty}\frac{4^{d}}{cN^{\textup{{val}}}B_{d}}\log\frac{6^{d}N^{\textup{{val}}}}{B_{d}\delta}=0$, and so this choice satisfies our earlier assumption. For this choice of $\tau$
@@
where in the final inequality we use that $B_{d}\leq 2^{d}$ since the unit ball is contained in the unit cube. For all $N^{\textup{{val}}}\geq\frac{\delta e^{c/2^{d}}}{3^{d}}$ the right-hand side is less than $\delta$; but $\delta<1,c\leq 1$ and so this holds for all $N^{\textup{{val}}}\geq 1$. ∎
D.3 Proof of Inconsistency of Existing Methods
@@ D.3.1 Inconsistency of the Holdout
See 4.1 We prove the strong result.
Proposition D.5 (Counterexample to the consistency of holdout).
Consider a single test point $(S,X,Y)$ satisfying Assumption 2.2. Suppose the test risk at such a point (Definition 2.3 with $M^{\textup{{test}}}=1$) is not constant as a function of $S$. Then there exists a test set (of size one) such that $\smash{\hat{R}_{\textup{{Hold}}}}$ is not a consistent estimator of the test risk on that test set.
@@
Because $e_{h}(S)$ is not a constant $\mathcal{S}$ contains at least two elements $S,S^{\prime}$ such that $e_{h}(S)\neq e_{h}(S^{\prime})$. Let $Q=\delta_{S}$ and $Q^{\prime}=\delta_{S^{\prime}}$. There there exists some $\gamma>0$ such that
$|R_{Q}(h)-R_{Q^{\prime}}(h)|=|e_{h}(S)-e_{h}(S^{\prime})|>2\gamma.$ (25)
If $\hat{R}_{\textup{{Hold}}}$ is not consistent for $R_{Q}(h)$, we are done. Otherwise, by the definition of consistency, for all $N\geq N_{0}$ with probability at least $1/2$,
$|(\hat{R}_{\textup{{Hold}}})^{(N)}(h)-R_{Q}(h)|<\gamma.$ (26)
where we use $(\hat{R}_{\textup{{Hold}}})^{(N)}$ to denote the estimator constructed using the first $N$ validation points, $D^{\textup{{val}}}_{N}$. By the reverse triangle inequality,
$|(\hat{R}_{\textup{{Hold}}})^{(N)}(h)-R_{Q^{\prime}}(h)|\geq|R_{Q}(h)-R_{Q^{\prime}}(h)|-|(\hat{R}_{\textup{{Hold}}})^{(N)}(h)-R_{Q^{\prime}}(h)|.$ (27)
Combining Equation 26 and Equation 27, for all $N\geq N_{0}$ with probability at least $1/2$
$|(\hat{R}_{\textup{{Hold}}})^{(N)}(h)-R_{Q^{\prime}}(h)|>2\gamma-\gamma=\gamma,$ (28)
which implies the holdout is not consistent for $R_{Q^{\prime}}(h)$. ∎
Proposition D.6.
Let $\ell(a,b)=\max(1,|a-b|)$. There exists a data-generating process, test set containing a single site and predictive method satisfying infill asymptotics such that the holdout converges in probability to 00, while $R_{Q^{\textup{{test}}}}=1$.
Proof.
Consider $\mathcal{S}=[0,1]$, $h\equiv 0$, $Y=S$, $S^{\textup{{test}}}=1$ and
$S^{\textup{{val}}}_{m}$ =$=\begin{cases}U_{m}&m\text{\,is prime},\\ 0&m\text{\,otherwise}.\end{cases}$ (29)
with $U_{m}$ independent and identically distributed uniform variables. By the infinitude of primes, there are infinitely many $m$ such that $S^{\textup{{val}}}_{m}$ is uniformly distributed and, for example by Proposition 3.3, this implies that $S^{\textup{{val}}}_{m}$ satisfies the infill assumption. On the other hand, by the prime number theorem, as $M^{\textup{{test}}}\to\infty$, the density of primes in the natural numbers tends to 00, and so
$\hat{R}_{\textup{{Hold}}}(h)=\frac{1}{M^{\textup{{test}}}}\sum_{m=1}^{M^{\textup{{test}}}}S^{\textup{{val}}}_{m}=\frac{1}{M^{\textup{{test}}}}\sum_{\begin{subarray}{c}m=1\\ \text{prime}\end{subarray}}^{M^{\textup{{test}}}}U_{m}\leq\frac{|\{m:1\leq m\leq M^{\textup{{test}}},m\text{\, prime}\}|}{M^{\textup{{test}}}}\to 0,$ (30)
and $R_{Q^{\textup{{test}}}}(h)=1$. ∎
D.3.2 Inconsistency of $1$-nearest neighbor Estimator
We now turn to $1$-nearest neighbor risk estimation and restate and prove Proposition 4.3: See 4.3 We again actually prove a strong result
Proposition D.7.
Assume any test point satisfies Assumption 2.2. Assume there exists a constant $c>0$ such that for any test point $(S,X,Y)$, $\mathrm{Var}[\ell(Y,h^{\chi}(S))|S]\geq c$. Next, consider an infinite sequence of validation sets as in Definition 3.2. Suppose there exists an $S^{\prime}\in\mathcal{S}$ such that for any $r>0$ and $N^{\textup{{val}}}>0$, |$|\{1\leq j\leq N^{\textup{{val}}}:d_{\mathcal{S}}(S^{\textup{{val}}}_{j},S^{\prime})=r\}|\leq 1$. Choose any $Q^{\textup{{test}}}$ such that $Q^{\textup{{test}}}(\{S^{\prime}\})>0$. Then there exists a $\delta\in(3/4,1)$ and a $C^{(\delta)}>0$ such that, for each $N^{\textup{{val}}}$, with probability at least $1-\delta$, $|R_{Q^{\textup{{test}}}}(h)\!-\!(\hat{R}_{\textup{{NN}},1})^{(N^{\textup{{val}}})}(h)|\geq C^{(\delta)}.$ Here $(\hat{R}_{\textup{{NN}},1})^{(N^{\textup{{val}}})}$ denotes the 1NN estimator associated to the first $N^{\textup{{val}}}$ data points and $\delta$. $C^{(\delta)}$ and $\delta$ do not depend on $N^{\textup{{val}}}$ or other properties of the sequence of validation data.
The technical condition |$|\{1\leq n\leq N^{\textup{{val}}}:d_{\mathcal{S}}(S^{\textup{{val}}}_{j},S^{\prime})=r\}|\leq 1$ ensures that the 1NN set for each point contains exactly $1$ point. That is, there are no ties. This condition would be satisfied for the infinite sequence of validation sets with probability one if, for instance, the validation points were chosen i.i.d. from a uniform measure on a compact set; cf. Proposition 3.3. Alternatively, it can be removed if any form of tie-breaking that selects a single nearest neighbor is used in defining the estimator.
@@ D.3.2 Inconsistency of $1$-nearest neighbor Estimator
D.3.3 Preliminary Result
We begin by proving a result that says that if a bounded random variable has a second moment bounded below by $C$, then it cannot be close to zero most of the time. We will apply this inequality to the second moment of the difference between the $1$-nearest neighbor estimator and the test risk to in our proof of Proposition 4.3.
Proposition D.8.
@@ D.3.3 Preliminary Result
Because $U\in[-A,A]$, $A^{2}-U^{2}$ is a non-negative random variable. Applying Markov’s inequality, for any $t>0$,
@@
Take $t=\frac{A^{2}-C}{\delta}$ which is greater than 00 because $\delta\in(1-\frac{C}{A^{2}},1)$. Then Equation 32 becomes,
@@
D.3.4 Proof of Inconsistency of $1$-nearest neighbor
We return to our proof of Proposition 4.3. The idea will be to consider a point prediction task and then apply Proposition D.8 to show that with some fixed probability, the $1$-nearest neighbor estimator is a fixed distance away from the test risk, even as $N^{\textup{{val}}}$ increases.
Proof of Proposition 4.3.
$\mathbb{E}|R_{Q^{\textup{{test}}}}(h)-(\hat{R}_{\textup{{NN}},1})^{(N^{\textup{{val}}})}(h)|^{2}\geq\mathbb{E}|(\hat{R}_{\textup{{NN}},1})^{(N^{\textup{{val}}})}(h)-\mathbb{E}[(\hat{R}_{\textup{{NN}},1})^{(N^{\textup{{val}}})}(h)]|^{2}.$ (37)
Because the $\epsilon^{\textup{val}}_{n}$ are independent the variance of $(\hat{R}_{\textup{{NN}},1})^{(N^{\textup{{val}}})}$ is additive
$\mathbb{E}|(\hat{R}_{\textup{{NN}},1})^{(N^{\textup{{val}}})}(h)-\mathbb{E}(\hat{R}_{\textup{{NN}},1})^{(N^{\textup{{val}}})}(h)|^{2}$ $=\sum_{n=1}^{N^{\textup{{val}}}}(w^{\textup{{NN}},\!1}_{n})^{2}\mathbb{E}[\ell(Y^{\textup{{val}}}_{n},h^{\chi}(S^{\textup{{val}}}_{n}))-e_{h}(S^{\textup{{val}}}_{n})]$ (38)
$\geq V\sum_{n=1}^{N^{\textup{{val}}}}(w^{\textup{{NN}},\!1}_{n})^{2}.$ (39)
where $40<V<\Delta^{2}/$ 4 is the assumed lower bound on the variance of $\ell(Y^{\textup{{val}}}_{n},h^{\chi}(S^{\textup{{val}}}_{n}))-e_{h}(S^{\textup{{val}}}_{n})$ and we have left implicit the dependence of the weights on $N^{\textup{{val}}}$. Also, |$|\{1\leq n\leq N^{\textup{{val}}}:d_{\mathcal{S}}(S^{\prime},S^{\textup{{val}}}_{n})=r|\leq 1$, implies $S^{\prime}$ has exactly one $1$-nearest neighbor in $S^{\textup{{val}}}_{1:N^{\textup{{val}}}}$, call the index of this neighbor $n(S^{\prime})$. Then
Combining Equation 39 and Equation 40
$\mathbb{E}[|R_{Q^{\textup{{test}}}}(h)-(\hat{R}_{\textup{{NN}},1})^{(N^{\textup{{val}}})}(h)|^{2}]\geq VQ^{\textup{{test}}}(\{S^{\prime}\})^{2}.$ (41)
We now apply Proposition D.8 with $U=|R_{Q^{\textup{{test}}}}(h)-(\hat{R}_{\textup{{NN}},1})^{(N^{\textup{{val}}})}(h)|$ to conclude that for $\delta\in(1-\frac{VQ^{\textup{{test}}}(\{S^{\prime}\})^{2}}{\Delta^{2}},1)$ with probability at least $1-\delta$
$|R_{Q^{\textup{{test}}}}(h)-(\hat{R}_{\textup{{NN}},1})^{(N^{\textup{{val}}})}(h)|\geq\Delta\sqrt{1-\frac{1-\frac{VQ^{\textup{{test}}}(\{S^{\prime}\})^{2}}{\Delta^{2}}}{\delta}}>0.$ (42)
As neither $\delta$ nor the right hand side of Equation 42 depend on $N^{\textup{{val}}}$, $1$-nearest neighbor is not consistent under infill asymptotics. ∎
Proposition D.9.
Consider $\mathcal{Y}=\{0,1\}$ and $\ell(a,b)=\begin{cases}0&a=b\\ 1&\text{otherwise}\end{cases}$, $\mathcal{S}=[0,1]^{d}$ and $S^{\textup{{val}}}_{m}\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}\mu$, where $\mu$ is any measure with density with respect to Lebesgue measure. Fix any $S\in\mathcal{S}$ and any predictive method $h$ and let $Q^{\textup{{test}}}=\delta_{S}$. Then,
$|\hat{R}_{\textup{{NN}},1}(h)-R_{Q^{\textup{{test}}}}(h)|\geq\min(\mathbb{E}[Y^{\textrm{test}}],1-\mathbb{E}[Y^{\textrm{test}}]).$ (43)
In particular, if $\mathbb{E}[Y^{\textrm{test}}]=1/2$, then one-nearest neighbor risk estimation has error $1/2$.